2010-12-02 (UT)
===============

Program: CL2 (YSOs, PI: Kraus et al.)

Observers: Stefan Kraus, PJ Goldfinger

B1  B2  B3
----------
S1  W1  E1
S2  W2  E2

(S1S2E1E2W1W2: 4-1-4-1-2-5)

CLIMB on beam 1,2,3

LT 18:00 = UT 02:00


*** S2E1W2, W2=ref.

S2 is wobbling a bit in the acquisition image,
one hour later it's better

LT 18:06 - HD15464 (fringe finding star)
   	   W2 tiptilt works OK for 3-4 minutes,
	   but then first the green points are moving out,
	   later also the white spots

	   realize that beam/telescope order was switched 
	   (E1 on B3 instead of B2), maybe this caused a misalignment? so we correct it.
	   problem with S2 time, restart telescope server
	   Laszlo is still working on S1, so we cannot switch to S1E1W1 yet
	   W2 is still drifting (TIPTILT SERVER says: "Failed to run message job.")

	   seeing 1-10 cm
	   B1B2=-0.199 cm, B2B3=0.789cm
	   

LT 20:02 - HD24365 (calibrator for RY Tau)
   	   E1 tip-tilt does not lock, increase NSUM to 5
	   recording data (750 Hz, NON-DEST, LONG)
	   B1B2=-0.209 cm, B2B3=0.758 cm
	   
LT 20:25 - RY Tau
	   B1B2=-0.209 cm, B2B3=0.758 cm
   	   recording data (750 Hz, NON-DEST, LONG)
	   most of the time fringe on B1B2
	   E1 tiptilt is working very bad (going in circles), so most of the time we don't have fringe in B2B3/B3B1, but at the end of data file, things improve
   	   B1B2=-0.208 cm, B2B3=0.757 cm

LT 20:46 - HD27848 (alternative calibrator for RY Tau)
   	   E1 tiptilt makes it worse than without tiptilt, so I use "GRAP" to center the star (works reasonably well)
	   B1B2=-0.218 cm, B2B3=0.762 cm
   	   recording data (750 Hz, NON-DEST, LONG)
	   (shutter files won't be representive of the photometry quite likely)

LT 21:03 - RY Tau
	   B1B2=-0.218 cm, B2B3=0.762 cm
   	   recording data (750 Hz, NON-DEST, LONG), better quality than last data set

LT 21:12 - HD32480
   	   B1B2=-0.217 cm, B2B3=0.754 cm
   	   recording data (750 Hz, NON-DEST, LONG), E1 in "GRAB" mode instead of tiptilt

LT 21:30 - RY Tau
   	   B1B2=-0.217 cm, B2B3=0.754 cm
   	   recording data (750 Hz, NON-DEST, LONG), E1 using tiptilt now (behaves better)
	   detector is saturating now (was no problem before, but better check again...!!!!)
   	   B1B2=-0.213 cm, B2B3=0.774 cm
	   recording data (750 Hz, DEST, LONG), E1 using tiptilt now
	   CLIMB GUI window freezes (CLIMB_1) during BEAM2 sequence, so I don't know whether the file has been saved
	   
LT 22:00 - HD32480, seeing says 1-6cm, but likely down-biased due to the strange tiptilt system performance
   	   B1B2=-0.213 cm, B2B3=0.774 cm
	   recording data (750 Hz, DEST, LONG), E1 using tiptilt now

LT 22:13 - HD32480
   	   B1B2=-0.213 cm, B2B3=0.774 cm
	   recording data (750 Hz, NON-DEST, LONG), E1 using tiptilt

LT 22:25 - SU Aur
   	   B1B2=-0.218 cm, B2B3=0.788 cm
	   recording data (750 Hz, NON-DEST, LONG), E1 using tiptilt
	   B1B2 baseline very clear in first 300 frames
	   sometimes slight saturation effects
	   found B2B3 later on, so take the later scans for all baselines
	   good quality (saturation not really an issue)

LT 22:41 - HD32480 (calibrator for SU Aur)
   	   B1B2=-0.222 cm, B2B3=0.787 cm
	   recording data (750 Hz, NON-DEST, LONG), E1 using tiptilt
	   
LT 22:52 - HD36724 (alternative calibrator for SU Aur, just in case something is funky with the standard calibrator)
   	   B1B2=-0.222 cm, B2B3=0.787 cm
	   recording data (750 Hz, NON-DEST, LONG), E1 using tiptilt

LT 23:05 - HD53510 (bright star for NIRO alignment)
   	   
LT 23:20 - HD47157 (calibrator for MWC147)
   	   NIRO crashed
   	   B1B2=-0.199 cm, B2B3=0.792 cm
	   recording data (750 Hz, NON-DEST, LONG), E1 no tiptilt, really bad injection!!!
	   E1 tiptilt does not work again, so B2B3 and B1B3 fringes are only very rarely present	   

LT 23:44 - MWC147
   	   B1B2=-0.199 cm, B2B3=0.792 cm
	   recording data (750 Hz, NON-DEST, LONG), E1 no tiptilt, really bad injection!!!
	   only fringes on B1B2 (S2W2 baseline), injection in E1 is too bad
   	   B1B2=-0.195 cm

LT 00:00 - HD52456 (calibrator for MWC147)
   	   B1B2=-0.199 cm, B2B3=0.792 cm
	   recording data (750 Hz, NON-DEST, LONG), E1 no tiptilt, really bad injection!!!
	   only fringes on B1B2 (S2W2 baseline)

LT 00:10 - changing configuration



*** S1E2W1, W1=B2=ref.

LT 00:20 - HD43386 (NIRO alignment, fringe finding star)

LT 00:3? - HD52456
   	   B1B2=-0.500 cm
	   cannot find B2B3 fringe, extensive searching, have to rehome carts
	   problem with E2 delay line (does not home to 0)
   	   B1B2=-0.500 cm, B2B3=+0.559 cm
	   recording data (750 Hz, DEST, LONG), interrupted because we want NON-DEST

LT 01:26 - HD47157, seeing 2-8 cm
	   recording data (1000 Hz, NON-DEST, LONG) (750Hz had 75 counts, so I go to 1000 Hz)
   	   B1B2=-0.509 cm, B2B3=+0.550 cm
	   recording data (750 Hz, DEST, LONG), 1000 Hz was better
   	   B1B2=-0.506 cm, B2B3=+0.550 cm

LT 01:50 - MWC147
   	   1000 Hz saturates, so I have to go to 750 Hz DEST
	   recording data (750 Hz, DEST, LONG)
	   cannot find fringes

LT 02:20 - HD47157
   	   go back to calibrator to find fringes again

LT 02:30 - HD50277
   	   finding fringes again	   
   	   B1B2=-0.499 cm, B2B3=+0.580 cm
	   recording data (750 Hz, DEST, LONG)
	   clouds towards the south - maybe this is part of the problem
   	   B1B2=-0.502 cm, B2B3=+0.581 cm

LT 02:40 - MWC147, seeing 4-5 cm
   	   S1 oscillates, but then it improves again
   	   recording data (750 Hz, DEST, LONG)
	   fringes not clear at all, but there might be something?!?
	   giving up

LT 02:54 - HD55137 (calibrator for MWC166 or HD58647)
   	   B1B2=-0.530 cm, B2B3=+0.534 cm
   	   recording data (1000 Hz, NON-DEST, LONG)

LT 03:05 - MWC166 (=HD53367)
   	   B1B2=-0.528 cm, B2B3=+0.546 cm
   	   recording data (1000 Hz, NON-DEST, LONG)
	   high contrast, very nice fringes
	   lost E2 during integration
	   piston seems quite bad

LT 03:20 - HD56405 (calibrator for MWC166)
   	   B2B3=+0.53 cm
	   cannot find B1B2 fringe?!?

LT 03:30 - HD55137 (calibrator for MWC166 or HD58647)
   	   B1B2=-0.531 cm, B2B3=+0.558 cm
   	   recording data (1000 Hz, NON-DEST, LONG)
	   it seems that the problem was that the system did not update the new star number

LT 03:45 - MWC166 (=HD53367)
   	   B1B2=-0.531 cm, B2B3=+0.558 cm
   	   recording data (1000 Hz, NON-DEST, LONG)
	   could record only lower frame number (130 frames) before running out of delay, but good quality!

LT 03:55 - HD65141 (calibrator for MWC166)
   	   B1B2=-0.551 cm, B2B3=+0.527 cm
   	   recording data (1000 Hz, NON-DEST, LONG)

	   
*** S2E2W2, W2=B2=ref.

LT 04:05 - HD53510 (bright star for setup)
   	   S2 oscillates very badly, so we cannot lock even on K=2 star.

LT 05:05 - HD73262 (calibrator for Beta CMi)
   	   S2 oscillated very badly, so we find mainly fringe on one baseline
   	   
LT 05:10 - Beta CMi
   	   S2 stopped oscillating
   	   
LT 05:15 - closing down